1)A particle is dropped from the top of the tower of height h. At the same instant another particle is projected vertically upwards from the bottom of the tower with such a velocity that it will be able to just reach the top of the tower. When and at what height from ground the two particles meet? 

Grade:

1 Answers

Sumit Majumdar IIT Delhi
8 years ago
Let both the particles meet after time t.
Hence,
The distance covered by the particle coming down from the top of the tower, would be equal to:
s=\frac{1}{2}gt^2
The distance covered by the particle going from ground would be given by:
h-s=ut-\frac{1}{2}gt^2
so, the time taken by both to meet, would be given by:
t=\frac{h}{u}
So the height from the ground would be equal to:
h-s=ut-\frac{1}{2}gt^2=u\times\frac{h}{u}-\frac{1}{2}g\times\left(\frac{h}{u} \right )^2=h-\frac{gh^2}{2u^2}